A.a particle having charge of 2.0×10^-4 is placed directly below and at a separation of 10 cm from the bob of a simple pendulum at rest .the mass of the bob is 100g .what charge should the bob be given si that the string becomes loose.

Grade:12th pass

1 Answers

Anirudh G Joshi
4 years ago
for the string to become loose, tension T in the string should become 0. This implies that downward force mg is balanced by the electrostatic repulsive force between the Bob and the charged  particle. Let the charge on the Bob be q.
F=mg
(9*10^9*2*10^-4*q)/(10*10^-2)^2=10*10^-3*9.8
18*10^(9-4+2)*q=10^-2*9.8
18*10^7*q=10^-2*9.8
q=10^-2*9.8/(18*10^7)
 therefore q=5.4*10^-10 C